# Agentar-Fin-R1: Enhancing Financial Intelligence through Domain Expertise, Training Efficiency, and Advanced Reasoning

### v1

@@ 4.5.1 Ablation Study on Label System and Weighted Training Framework
To evaluate the effectiveness of our proposed label-guided weighted training framework, we conduct an ablation study under constrained data regimes. The primary goal is to demonstrate that, by leveraging a structured task label system and difficulty-aware sample weighting, the model can achieve comparable or even superior performance with significantly fewer training samples.
We compare the following training configurations across different data budget constraints, all of which employ the SFT training method:
@@
All configurations are evaluated on the same downstream tasks, using the benchmark datasets: Fineval, Fineva, FinanceIQ, Finova, MATH, and GPQA. We report task-specific accuracy and overall average performance. The model used for the experiment is Qwen3-8B.
Training Strategy Financial Datasets General Reasoning Average
Fineval Fineva FinanceIQ Finova MATH GPQA All Datasets
@@
Table 4: Performance comparison across different training strategies under constrained and full data budgets. We evaluate our method with 10%, 30%, and 50% of the training data to demonstrate its effectiveness across various data budget constraints. Datasets are grouped into Financial and General Reasoning. Bold indicates the best performance among all configurations.
@@
  • Even with only 10% of the training data (30k samples), our approach achieves competitive performance (76.68 average), highlighting the efficiency of the label-guided weighted training framework
  • The performance progressively improves as we increase the data budget: 10% → 30% → 50% (76.68 → 77.34 → 78.10)
  • Our method consistently outperforms both label-only stratified sampling and random sampling baselines across all evaluation datasets
@@
  • Label System Impact: Label-only stratified sampling (76.75) outperforms random sampling (75.08) by 1.67, demonstrating the value of structured sampling
  • Weighting Mechanism Impact: Our full framework (78.10) further improves upon label-only sampling by 1.35, validating the effectiveness of difficulty-aware weighting
  • Combined Effect: The synergy between label system and weighting mechanism provides a total improvement of 3.02 over random sampling
@@ 4.5.2 Ablation Study on Two-Stage Training Pipeline
All configurations are evaluated on the same benchmark datasets: Fineval, Fineva, FinanceIQ, Finova, MATH, and GPQA.
Training Strategy Financial Datasets General Reasoning Average
Fineval Fineva FinanceIQ Finova MATH GPQA All Datasets
@@
Table 5: Performance comparison across different training strategies. Datasets are grouped into Financial and General Reasoning categories. Bold indicates the best performance among all training configurations.
@@ 5 Conclusion
Training Efficiency: We propose a novel weighted training framework incorporating difficulty-aware sample estimation and tow-stage optimization strategies that systematically maximizes data utilization efficiency while achieving superior convergence properties. This methodology enables targeted optimization across heterogeneous financial task distributions without incurring the computational penalties typically associated with domain specialization, thereby establishing a new paradigm for efficient domain adaptation. Our sophisticated financial label system further enhances training efficiency by enabling strategic sample selection and curriculum learning approaches that optimize resource allocation.

### v2

@@ 4.5.1 Ablation Study on Label System and Weighted Training Framework
To evaluate the effectiveness of our proposed label-guided weighted training framework, we conducted an ablation study under constrained data regimes. The primary goal is to demonstrate that, by leveraging a structured task label system and difficulty-aware sample weighting, the model can achieve comparable or even superior performance with significantly fewer training samples.
We compare the following training configurations across different data budget constraints, all of which employ the SFT training method.
@@
All configurations are evaluated on the same downstream tasks, using the benchmark datasets: Fineva, FinEval, FinanceIQ, Finova, MATH-500, and GPQA-diamond. We report task-specific accuracy and overall average performance. The model used for the experiment is Qwen3-8B.
Training Strategy Financial General Average
Fineva FinEval FinanceIQ Finova MATH GPQA All Datasets
@@
Table 4: Performance comparison across different training strategies under constrained and full data budgets. We evaluate our method with 10%, 30%, and 50% of the training data to demonstrate its effectiveness across various data budget constraints. Datasets are grouped into Financial and General. Note that MATH refers to MATH-500 and GPQA refers to GPQA-diamond. Bold indicates the best performance among all configurations.
@@
  • Even with only 10% of the training data (30k samples), our approach achieves competitive performance (76.68 average), highlighting the efficiency of the label-guided weighted training framework.
  • The performance progressively improves as we increase the data budget: 10% → 30% → 50% (76.68 → 77.34 → 78.10).
  • Our method consistently outperforms both label-only stratified sampling and random sampling baselines across all evaluation datasets.
@@
  • Label System Impact: Label-only stratified sampling (76.75) outperforms random sampling (75.08) by 1.67, demonstrating the value of structured sampling.
  • Weighting Mechanism Impact: Our full framework (78.10) further improves upon label-only sampling by 1.35, validating the effectiveness of difficulty-aware weighting.
  • Combined Effect: The synergy between label system and weighting mechanism provides a total improvement of 3.02 over random sampling.
@@ 4.5.2 Ablation Study on Two-Stage Training Pipeline
All configurations are evaluated on the same benchmark datasets: Fineva, FinEval, FinanceIQ, Finova, MATH-500, and GPQA-diamond.
Training Strategy Financial General Average
Fineva FinEval FinanceIQ Finova MATH GPQA All Datasets
@@
Table 5: Performance comparison across different training strategies. Datasets are grouped into Financial and General categories. Note that MATH refers to MATH-500 and GPQA refers to GPQA-diamond. Bold indicates the best performance among all training configurations.
@@ 5 Conclusion
Training Efficiency: We propose a novel weighted training framework incorporating difficulty-aware sample estimation and two-stage optimization strategies that systematically maximizes data utilization efficiency while achieving superior convergence properties. This methodology enables targeted optimization across heterogeneous financial task distributions without incurring the computational penalties typically associated with domain specialization, thereby establishing a new paradigm for efficient domain adaptation. Our sophisticated financial label system further enhances training efficiency by enabling strategic sample selection and curriculum learning approaches that optimize resource allocation.